he following information is available regarding the total manufacturing overhead of Olaf Company for a recent four-month period.

Machine Hours Mfg. Overhead
April 78,000 $ 180,000
May 50,000 $ 163,000
June 109,000 $ 251,500
July 105,000 $ 192,000

Using the high-low method, compute the variable element of manufacturing overhead cost per machine hour.

Using the high-low method, compute the fixed element of Olaf's monthly overhead cost.

Olaf's projected August operations will require approximately 160,000 machine hours. Using the high-low method, compute total manufacturing overhead estimated for August.
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(251500-163,000)/(109,000-50000)

=$1.5 per machine hour

Total fixed cost=251500-(1.5*109,000)

=$88000

Total cost for August=Total variable cost+Total fixed cost

=(1.5*160,000)+88000

=$328000